# Iodine and Bromine Radical Reactions in Atmospheric Mercury Oxidation

### v1

@@ I Introduction
In this paper we use computational methods to investigate the chemical interactions of mercury with bromine and iodine in the atmosphere, focusing in particular on the areas around the poles and Arctic Mercury Depletion Events (AMDEs). Interest in halogens reacting with mercury is driven by their role in promoting the oxidation of Hg to Hg(II) during AMDEs, leading to mercury depostion to the surface. Research focus has mainly been on the reactions between Br and Hg.[1, 2, 3, 4] It has been suggested that iodine could also play an important role, as an oxidant or by acting to increase the concentration of Br.[5, 6] Due to increasing atmospheric iodine concentrations in the Arctic[7], here we will study the effects of this increase on mercury oxidation and contrast the iodine chemistry with the bromine chemistry. This will be done using computational chemistry to examine the rates of mercury reactions with iodine.
Mercury persists in the atmosphere for about six to eighteen months, which is long enough for it to spread all over the world.[8] Over time, mercury pollution from the populated areas of the Northern Hemisphere reaches the Arctic. In springtime high levels of halogen radicals can be generated photochemically, and tropospheric ozone layer in the Arctic is temporarily depleted and Hg(0) is oxidised.[9] Mercury deposition (and aqueous solubility) increases when it is oxidised, enabling mercury to enter biological ecosystems and rendering them hazardous to humans and animals [4]. Currently, the level of mercury found in the blood of children in areas of the Arctic region is sufficient to cause nerve damage.[10, 11] The mercury level at present is 500 % above its natural atmospheric level [12].
@@ I.1 The atmospheric chemistry of mercury
[…]
Globally the main oxidants for Hg(0) are ozone (O), generated photochemically [18, 19], hydroxyl radicals, nitrogen oxides (NO) and hydrogen peroxide (HO).[4] The nitrate radical (NO) also plays a role at night.[19] While these are the main reactants, in the Arctic atmosphere during springtime, halogen reactions with unoxidized mercury prove to be more significant than the reactions described above. This is due to emissions of halogen compounds and radicals into the atmosphere and at the atmospheric boundary layer of seawater, snow, and ice. The halogen reactions occur mainly in sea salt aerosols, particularly in coastal areas.[4, 19] Oxidized mercury is observed in the atmosphere in low concentrations because it readily dissolves into wet surfaces due to its low vapor pressure.[20, 18, 21, 22]
@@ I.2 Halogen reactions in the atmosphere
The production of reactive halogen species radically changes chemical processes in the Arctic atmospheric boundary layer upon the return of sunlight in the polar spring.[23, 9] A rapid drop in ozone from 30-40 ppb down to 10 ppb has been observed during springtime ozone depletion events in the Arctic.[24, 25] A correlation between bromine and ozone suggested the catalytic destruction of ozone by bromine atoms, [26] but modeling suggests this reaction is not fully understood, and other halogens’ reactivities are not adequately known.[27] Intriguingly, Spolaor et al. found that Hg and I are correlated with each other and with sunlight, while Br did not show a diurnal cycle.[28] Initial observations of inorganic bromine (Br, BrO, HOBr), [23] in the polar region have later been updated with the findings of inorganic chlorine (Cl, ClO)[29, 30] and iodine (IO, HIO).[31, 32] Even if molecular iodine (I) has not been previously observed in the Arctic, observations have been made in midlatitude marine and coastal sites,[7] along the Artic and Antarctic coast.[33] Furthermore, IO has been observed in the Antarctic and the sub-Arctic.[34]
@@
Globally, inorganic iodine compounds released from the ocean account for roughly 3/4 of atmospheric iodine. The remaining atmospheric iodine is estimated to come from organoiodine species (e.g., CHI, CHI, etc.).[35, 36] It has been demonstrated in laboratory studies that oceanic emission of hypoiodous acid (HOI) and molecular iodine (I) follow the deposition of tropospheric ozone (O) on the ocean surface, and therefore including reactions with iodide (I) ions.[37, 38] The increase in global ocean iodine emissions may be in response to a human-driven increase in the level of tropospheric ozone during the industrial period and/or due to increasing Arctic sub-ice biological activity in response to climate change.[39, 40]
I.3 Ice core measurements prove rising iodine levels
As described by Saiz-Lopez et al.,[7] the role of iodine chemistry may increase as the Arctic warms, given the prevalence of iodine chemistry in the marine midlatitudes. Studies of ice cores from Renland on Greenland show that iodine concentrations increased from 1950 to 2010, after being fairly stable from 1750 to 1950. Cuevas[40] attributes the increase of iodine in the ice cores and the atmosphere to human influences - increased tropospheric ozone and recent Arctic warming - leading to an increase in Arctic sub-ice biological activity. An increase in atmospheric iodine levels followed by deposition to ice-covered areas would make local heterogeneous processing and recycling of iodine in ice/snow more efficient, a mechanism supported by recent measurements in the Arctic.[41, 42] Remarkably, from 1990 to 2011, iodine concentrations increased despite stable atmospheric ozone concentrations.[40] Mole fractions (IO) of 1.5 ppt were observed at Alert, Canada in 2015 [31]. Particle formation arising from ppt () levels of reactive iodine species such as iodic acid (HIO) has been observed in the Arctic.[32] These observations, combined with the difficulty of performing laboratory experiments and theoretical investigations with these elements, underscore the need for additional work to characterize iodine’s impact on mercury deposition.[34]
@@ I.3 Ice core measurements prove rising iodine levels
There is also circumstantial evidence suggestive of iodine’s role. In the Arctic saline snowpack, photochemical reactions have been demonstrated involving Br, Cl and BrCl.[43, 44, 30] Photochemical reactions with these species have also been shown in laboratory experiments.[45, 46, 47, 48, 49, 50, 51, 52] Further, it has been shown that I and triiodide (I ) can be photochemically produced in Antarctic snow spiked with iodide (I) (1–1,000 M).[42] It has also been shown that iodate (IO) can be photochemically active in frozen solutions.[53] The referenced studies suggest the probability of photochemical production of I is similar to that of Br, Cl, and BrCl in the Arctic snowpack.[43, 30, 34] Therefore, this article will aim to investigate rates of atmospheric reactions between mercury and iodine and help to estimate the impact of rising atmospheric iodine concentrations over the arctic.
@@
In this article, we investigate reactions R1-R3 described below. The reactions are suggested by Schroeder et al.[18] and Jiao and Dibble.[2]
Hg + M +X HgX |
HgX + Y YHgX |
HgX + Y HgXY Hg + XY |
@@
Here, M is a third body collision partner (N, O..), and X = Br, I, and Y = BrO, ClO, IO, NO, HO, I, or Br. The two last reactions were proposed by Jiao and Dibble as likely propagators in the mechanism. In this article, we focus on reaction chains starting from reaction R1 where X = I and Br, and the results will be compared with results of X = Br from other calculations and from literature where this is possible. The goal is to evaluate the extent to which increasing iodine levels in the Arctic will influence AMDEs.
@@ II.1 Computational Details
In this study we compute reaction energies and reaction rates using the methods of computational quantum chemistry.
The CCSD(T) method with effective core potential (ECP) and a triple zeta basis set is a well-established computational method for calculating reaction energies and properties of mercury at equilibrium geometries.[54, 55] For this project, the specific method used is CCSD(T) with the minimally augmented basis set ma-def2-TZVP, with effective core potentials for relativistic treatment.[56, 54, 55] Previous theoretical investigations of mercury compounds have shown that this method is good for both geometry optimizations and frequency calculations and has been used for accurate calculations on mercury compounds. This is due to the large contribution from the correlation energy compared to that of the contribution of relativistic effects to the energy, though the latter is still significant.
Rate coefficients have been calculated using variational transition state theory.[57] Relaxed geometry scans of the dissociation reactions were performed using CASSCF and the same ma-def2-TZVP basis set to obtain energies and frequencies at intermediate geometries, as input for the variational transition state theory. Preliminary calculations at the unrestricted CCSD(T) level gave many spurious results at intermediate geometries due to spin-contamination, forcing us to employ a proper multi-configurational approach. CASPT2 single-point energy calculations were performed on the geometries, again with the ECP - ma-def2-TZVP basis set. Frequency calculations were performed using either the CASPT2 or CASSCF method with ZORA for the relativistic approximation and ZORA-TZVP basis set. It was necessary to use the ZORA relativistic approximation since the ECP approximation gave results that gave many imaginary frequencies. About 47 intermediate single point energy calculations with the CASPT2/ma-def2-TZVP level of theory and frequency calculations were carried out on the same geometries with the CASPT2 or CASSCF method and ZORA-TZVP basis set. All electronic structure calculations were carried out with the ORCA program.[58, 59, 60] Input files for the calculations in Orca and Ktools can be found in the supplementary Information Section 3.1.
@@ II.2 Kinetics
The formulation used in this study is based on the one used in the Multiwell program package.[61, 62] All relevant equations are detailed in the manual, primarily in the sections describing the theory behind Ktools (p. 114) and sctst (p. 209) subprograms. Equation (1) outlines the method for calculating rate coefficients in Ktools for uni-molecular dissociation reactions, averaged over the electronic, vibrational and rotational energy at a specific temperature.
@@ III.1.1 Reaction R1 with X = Br or I
The first reaction step, creating the Hg(I) intermediate with iodine or bromine, is similar for the two reactions and is strightforward to compare. The only geometric difference is the bond length in the HgX radicals that differs by 8% (0.22 Å). This also results in a change in frequency of 1.12 cm.
There is also a difference in the reaction of Gibbs free energy between the reaction of Hg+I (-5.17 kcal/mol) and Hg+Br (-10.52 kcal/mol). The difference between the two reactions is thus -5.35 kcal/mol. This is an important difference since it would favor the formation of HgBr radicals over HgI radicals. Geometries and frequencies can be found in the supplementary material section 1.1 Table S1.
III.1.2 Reaction R2 with X = Br or I and Y = Br, I, BrO, ClO, IO, or HO
The difference in bond length for HgI and HgBr is 7% (0.19 Å). The XHgY molecules have similar geometries (X=halogen; and Y = BrO, ClO, IO, and HO). The difference in bond length for the X-HgOX bond shrinks by about 7% (0.17 Å) when interchanging iodine with bromine in the XHgOX molecules.
Furthermore, there is a general increase of the X-HgOX and XHgO-X bond lengths with an increase of the XHgO{X} halogen and a decrease in bond length for the XHg-OX bond. The increase of atomic number of the same halogen also increases the HgOX bond angle, but the XHgO bond angle changes less than a degree.
The largest difference in bond angles between any of these Hg(II)-products is 3.5 degrees between BrHgOCl and BrHgOI for the XHg-O-X bond angle. The largest difference in bond length is 14% (0.28 Å) between IHgOI and IHgOCl for the XHgO-X bond. In the case of the XHgOOH complexes, IHgOOH and BrHgOOH are very similar in geometry, only differing noticeably in the bond length of the X-HgOOH bond (7% or 0.17 Å). See section 1.3 in the supplementary material Table S5 - S8, S31 - S32.
For XHgOX and HgX the largest energy difference between reactant and product is for the HgBr + Br HgBr reaction with a reaction Gibbs free energy change of -61.09 kcal/mol. The stabilization by the Gibbs free energy change decreases from XHgOI over XHgOBr to XHgOCl and XHgOOH. Here the reaction Gibbs free energy is always less negative for the IHgOX molecules than the BrHgOX molecules. The smallest reaction Gibbs free energy is -25.15 kcal/mol in the case of the IHgOOH reaction. There is an energy difference of between 2.65 and 2.70 kcal/mol when interchanging the Br and I at the XHg radical in reaction R2. There is a general increase in reaction Gibbs free energy with increasing atomic number of the halogen in OX radical of reaction R2. The reaction Gibbs energy of the formation of BrHgOOH is 2.983 kcal/mol larger than the Gibbs free energy for the reaction producing IHgOOH. The results for reaction R2 can be seen in Table 1.
@@ III.1.2 Reaction R2 with X = Br or I and Y = Br, I, BrO, ClO, IO, or HO
III.1.3 Reaction R2 with X = Br or I and Y = NO
The X-HgNO reaction path exhibits two intermediates and two transition states.
@@ III.1.3 Reaction R2 with X = Br or I and Y = NO
These can be understood as an example of reaction R2. In the first step of this reaction pathway, the loan pair in nitrogen binds to mercury and creates the intermediate XHgNO. Thereafter, a transition state (TS1) creates a barrier for reaching the anti-conformer XHgONO-anti. Another transition state (TS2) creates a barrier for reaching the syn-conformer XHgONO-syn, see Figure 1.
BrHgNO
”TS1”
@@
Geometries for products and intermediates have been compared to the result from Jiao and Dibble.[56] In Jiao and Dibble’s article, the geometries of stationary points (reactants, products, and well-defined transition states) and points along the minimum energy path for reactions were optimized using the PBE0 density functional with the aug-cc-pVTZ basis set and relativistic effective core potentials and then verified by harmonic vibrational frequency analyses at the same level of theory. The energies of stationary points were refined using the CCSD(T) method in conjunction with the aug-cc-pVTZ basis set. In the present study, the geometries were re-optimizated and harmonic frequencies were calculated with the CCSD(T) method, ECPs and the ma-def2-TZVP basis set.
We investigate (TS1) from Jiao and Dibble,[56] by re-optimizing the PBE0 geometry of the transition state at the CCSD(T) level of theory; only a local minimum is found. The CCSD(T) geometry of this local minimum has a difference of 6.50 for the Br-Hg-NOO bond angle, 5.30 for the XHgO-N-O bon angle, 40.30 for the XHg-N-O(1)O(2) and 35.00 for the XHg-N-O(2)O(1).
In this article, the largest difference in bond length of the same configuration is for TS1, where the difference is 0.25 Å for the XHg-ONO bond and the largest difference in bond angle is 0.9 (XHgO-N-O) for the XHgONO-syn conformers. Further information can be found in the supplementary material section 1.5 Tables S16 - S30.
The relative energies for the reaction’s IHgNO products are similar to those of BrHgNO, because the geometries along the reaction path are very similar, see Table 2. Due to the difference in the nature of the local minimum ”TS1” for BrHgONO in the present work compared to Jiao and Dibble,[56] the difference in the Gibbs free energy and the enthalpy for ”TS1” is also different. When using Jiao’s PBE0 geometry as a basis for a frequency calculation the energy difference between TS1 and the reactant for BrHgONO is 13.75 kcal/mol (298 K). However, when creating a new transition state optimization at the CCSD(T) level, the energy difference is only 1.38 kcal/mol (298 K). TS1 for the IHgONO has an energy difference of 2.47 kcal/mol (298 K) when optimized with the CCSD(T) method, though the same problem applies to IHgONO ”TS1”. The highest energy barrier along the reaction path for the BrHgNO and IHgNO complex is TS2 with an energy barrier of 5.93 kcal/mol for BrHgNO and 7.12 kcal/mol for the iodine complex.

### v2

# Bromine and Iodine in Atmospheric Mercury Oxidation

@@ I Introduction
In this paper we use computational methods to investigate the chemical interactions of mercury with bromine and iodine in the atmosphere, focusing in particular on local phenomina and Arctic Mercury Depletion Events (AMDEs). Interest in halogens reacting with mercury is driven by their role in promoting the oxidation of Hg to Hg(II) during AMDEs, leading to mercury depostion to the surface. Research focus has mainly been on the reactions between Br and Hg. [22, 36, 32] It has been suggested that iodine could also play an important role, as an oxidant or by acting to increase the concentration of Br. [12, 2] Due to increasing atmospheric iodine concentrations in the Arctic [58], we have studied the effects of this increase on mercury oxidation and contrast the iodine chemistry with the bromine chemistry. This will be done using computational chemistry to examine the rates of mercury reactions with iodine.
Mercury persists in the environment long enough for it to be transported all over the globe. [74] Over time, mercury pollution from populated areas of the Northern Hemisphere reaches the Arctic. In springtime high levels of halogen radicals can be generated photochemically, and tropospheric ozone in the Arctic is temporarily depleted, and Hg(0) is oxidised. [73] Mercury deposition (and aqueous solubility) increases when it is oxidised, enabling mercury to enter biological ecosystems and rendering them hazardous to humans and animals [27].
@@ I.1 The atmospheric chemistry of mercury
When entering the atmosphere, the residence time of elemental unoxidized mercury is between 4 and 7 months, which is enough for it to scatter across the Northern Hemisphere. [60, 21, 33, 31] Oxidized mercury, on the other hand, is soluble in water, and its residence time is between hours and days. [63, 2] If the temperature increases by 10 °C, the amount of evaporated Hg(0) doubles [38] linking mercury pollution with human activity and climate change. [38, 83, 34] Mercury can undergo multiple deposition and evaporation cycles furthering it’s spread.
Globally the main oxidants for Hg(0) are OH and Br radicals, and the main reactants for the further reaction from Hg(I) to Hg(II) are O3, NO2, and OOH. [65] Shah et al. estimate that Br radicals are responsible for about 75% of the oxidation of Hg(0) globally and a lifetime of mercury in the atmosphere of 5.5 months. While these are the main reactants, in the Arctic atmosphere during springtime, halogen reactions with unoxidized mercury prove to be even more significant. This is due to emission of halogen compounds and radicals into the atmosphere including from seawater, snow, and ice. [64, 61, 42, 43]
In the troposphere, it is estimated that gas phase Hg(II) is the most abundant form, comprising 78% of Hg(II) mass, while Hg(II) processed by aerosol and cloud droplets comprises 22%. The total deposition flux of Hg(II) is 5.5 Gg/a, where 71% of the Hg(II) disposition is situated over the oceans. [65]
A main argument put suggested in previous studies investigating the reaction between Hg and I focus on the low thermodynamic stability of the Hg(I) complex, leading to rapid decomposition and therefore not contributing to mercury depletion. [26, 17, 4] A similar rationale has been proposed for the reaction between Hg and OH, with similar thermodynamic stabilization energy, [13] though Dibble et al. find that while the Hg + OH pathway contributes <1% of the conversion of Hg(0) to Hg(II) globally, it contributes up to 83% in polluted regions, [23] thereby underscoring the complexity of mercury’s atmospheric chemistry and its sensitivity to environmental conditions. Though we do not investigate the previous rates calculated for Hg + I HgI, a similar argument as for the reaction between Hg(I) and OH (D0 = 11.02 kcal/mol) could be put forward, as a reaction preceding the reactions of IHg (D0 = 10.03 kcal/mol) with other atmospheric species.
@@ I.2 Halogen Reactions in the Atmosphere
Reactive halogen species significantly alter chemical processes in the Arctic atmospheric boundary layer during polar spring. [68, 73] Observations have shown rapid ozone depletion events, with ozone levels dropping from 355 ppb to 10 ppb. [9, 48] This has been linked to catalytic ozone destruction by halogen atoms. [10, 35, 79] Notably, studies have found correlations between mercury (Hg) and iodine (I), while bromine (Br) did not show a diurnal cycle. [72] Although molecular iodine (I2) has not been directly observed in the Arctic, iodine species like IO have been detected in the Antarctic and sub-Arctic regions. [3, 56]
@@
Approximately 75% of atmospheric iodine originates from inorganic compounds released from the ocean, with the remainder from organoiodine species. [14, 52] Laboratory studies show that oceanic emissions of hypoiodous acid (HOI) and molecular iodine (I2) are linked to the deposition of tropospheric ozone on the ocean surface, involving reactions with iodide (I-) ions. [15, 41] This suggests that increased tropospheric ozone and Arctic warming may enhance iodine emissions. [51, 18]
I.3 Ice core measurements indicate rising iodine levels
Recent studies indicate that iodine chemistry may become more significant as the Arctic warms. [58] Ice core analyses from Greenland show a marked increase in iodine concentrations since 1950, likely due to human influences such as increased tropospheric ozone and Arctic warming. [18] This rise in atmospheric iodine enhances local processing and recycling of iodine in ice and snow. [59, 40] Notably, mole fractions of IO up to 1.5 ppt were observed at Alert, Canada. [87] Particle formation from ppt levels of reactive iodine species like iodic acid (HIO3) has been detected in the Arctic. [69] These findings highlight the need to investigate the rates of atmospheric reactions between mercury and iodine to assess the impact of rising iodine levels.
@@ I.3 Ice core measurements indicate rising iodine levels
Photochemical reactions in the Arctic saline snowpack involving halogens like Br2, Cl2, and BrCl have been demonstrated both in the field and laboratory. [53, 20] Similar photochemical production of I2 and triiodide (I) from iodide has been shown in Antarctic snow. [40] Additionally, iodate (IO) can be photochemically active in frozen solutions. [25] These studies suggest that the photochemical behavior of iodine may be comparable to that of bromine and chlorine in the Arctic. Therefore, investigating the rates of atmospheric reactions between mercury and iodine is crucial to estimate the impact of rising atmospheric iodine concentrations.
@@
In this article, we investigate reactions R1-R3 described below. The reactions are suggested by Goodsite, [26] Schroeder et al. [61] and Jiao and Dibble. [36]
| Hg + M +X HgX + M | (R1) | |
| HgX + Y + M YHgX +M | (R2) | |
| HgX + Y + M HgXY +M Hg + XY + M | (R3) |
@@
Here, M is a third body collision partner (N2, O2..), and X = Br, I, and Y = BrO, ClO, IO, NO2, HO2, I, or Br. The two last reactions were proposed by Jiao and Dibble as likely propagators in the mechanism. In this article, we focus on reaction chains starting from reaction R1 where X = I and Br. The results will be compared with results of X = Br from other calculations and from literature when possible. The goal is to produce thermodynamic and kinetic data so evaluation of the increasing iodine levels local influence on AMDEs can be made possible in the future.
@@ II.1 Computational Details
Reaction energies and rate coefficients were computed using computational quantum chemistry methods. The CCSD(T) method with effective core potentials (ECPs) and the minimally augmented basis set ma-def2-TZVP was employed for calculating reaction energies and properties of mercury at equilibrium geometries, incorporating relativistic effects [39, 5, 37]. These previous theoretical investigations of mercury compounds have shown that this method is good for both geometry optimizations and frequency calculations and has been used for accurate calculations on mercury compounds. This is due to the large contribution from the correlation energy compared to that of the contribution of relativistic effects to the energy, though the latter is still significant.
Rate coefficients were calculated using variational transition state theory (VTST) [78]. Relaxed geometry scans of the dissociation reactions were performed using the CASSCF method with the ma-def2-TZVP basis set to obtain energies and frequencies at intermediate geometries as input for VTST. Due to spin-contamination issues with unrestricted CCSD(T) at intermediate geometries, a multiconfigurational approach was necessary.
Single-point energy calculations were conducted using CASPT2 with the same basis set. Frequency calculations employed either CASPT2 or CASSCF methods with the ZORA relativistic approximation and the ZORA-TZVP basis set, as the ECP approximation resulted in multiple imaginary frequencies. Approximately 47 intermediate single-point energy and frequency calculations were performed. See Figure 2
@@ II.2 Kinetics
The formulation used in this study is based on the one used in the Multiwell program package. [8, 86] All relevant equations are detailed in the manual, primarily in the sections describing the theory behind the Ktools (p. 114) and sctst (p. 209) subprograms. Equation (1) outlines the method for calculating rate coefficients in Ktools for uni-molecular dissociation reactions, averaged over the electronic, vibrational and rotational energy at a specific temperature.
@@ III.1.1 Reaction R1 with X = Br or I
There is a difference in Gibbs free energy of the reaction between the reaction of Hg+I (-5.17 kcal/mol) and Hg+Br (-10.52 kcal/mol). The difference between the two reactions is thus -5.35 kcal/mol. This is an important difference since it would favor the formation of HgBr radicals over HgI radicals. Geometries and frequencies can be found in the supplementary material section 1.1 Table S1.
III.1.2 Reaction R2 with X = Br or I, and OY= BrO, ClO, IO, or HO2
For XHgY and HgX2 the largest energy difference between reactant and product is for the HgBr + Br HgBr2 reaction with a reaction Gibbs free energy change of -61.09 kcal/mol. The stabilization by the Gibbs free energy change decreases from XHgOI over XHgOBr to XHgOCl and XHgOOH. Here the reaction Gibbs free energy is always less negative for the IHgOY molecules than the BrHgOY molecules. The smallest reaction Gibbs free energy is -25.15 kcal/mol in the case of the IHgOOH reaction. There is an energy difference of between 2.65 and 2.70 kcal/mol when interchanging the Br and I at the XHg radical in reaction R2. There is a general increase in reaction Gibbs free energy with increasing atomic number of the halogen in OY radical of reaction R2. The reaction Gibbs energy of the formation of BrHgOOH is 3.0 kcal/mol larger than the Gibbs free energy for the reaction producing IHgOOH. The results for reaction R2 can be seen in Table 1.
@@ III.1.2 Reaction R2 with X = Br or I, and OY= BrO, ClO, IO, or HO2
III.1.3 Reaction R2 with X = Br or I and Y = NO2
The X-HgNO2 reaction path exhibits two intermediates and two transition states.
@@ III.1.3 Reaction R2 with X = Br or I and Y = NO2
These can be understood as an example of reaction R2. In the first step of this reaction pathway, the lone pair in nitrogen binds to mercury and creates the intermediate XHgNO2. Thereafter, a transition state (TS1) creates a barrier for reaching the anti-conformer XHgONO-anti. Another transition state (TS2) creates a barrier for reaching the syn-conformer XHgONO-syn, see Figure 1.
BrHgNO2
TS1
@@
Geometries for products and intermediates have been compared to the result from Jiao and Dibble. [37] In Jiao and Dibble’s article, the geometries of stationary points (reactants, products, and well-defined transition states) and points along the minimum energy path for reactions were optimized using the PBE0 density functional with the aug-cc-pVTZ basis set and relativistic effective core potentials and then verified by harmonic vibrational frequency analyses at the same level of theory. The energies of stationary points were refined using the CCSD(T) method in conjunction with the aug-cc-pVTZ basis set. In the present study, the geometries were re-optimized, and harmonic frequencies were calculated with the CCSD(T) method, ECPs, and the ma-def2-TZVP basis set.
In this article, the largest difference in bond length of the same configuration is for the halogen mercury bond X-HgONO with a change in bond distance between 0.16 Å - 0.18 Å and the largest difference in bond angle is 0.9∘ (XHgO-N-O) for the XHgONO-syn conformers. Further information can be found in the supplementary material section 1.5.
The relative energies for the reaction’s IHgNO2 products are similar to those of BrHgNO2, because the geometries along the reaction path are very similar, see Table 2. At TS1 for BrHgONO the energy difference is 12.88 kcal/mol (298 K). TS1 for the IHgONO has an energy difference of 12.97 kcal/mol (298 K) when optimized with the CCSD(T) method. The TS2 energy barrier for the BrHgNO2 and IHgNO2 complex is 5.93 kcal/mol for BrHgNO2 and 7.12 kcal/mol for the iodine complex.